If an R = 1-kΩ resistor, a C = 1-µF capacitor, and an L = 0.2-H inductor are connected in
series with a V = 150 sin (377t) volts source, what is the maximum current delivered by the source?
1
Expert's answer
2020-11-09T14:18:53-0500

Solution: Based on the datas given in the question a circuit diagram can be constructed as shown in the below figure.





From this we can calculate capacitive reactance = "X_c = 1/C\omega = 1/(10^{-6} \times377) = 2652.52\varOmega"


The inductive reactance =

"X_L = L\omega = 0.2\times377 = 75.4 \varOmega."


The rms value of input voltage "= 150/\sqrt{2} = 106.07V"


Hence an equivalent circuit in the frequency domain incorporating the sign of capacitive and inductive reactance is as shown below





Total impedance of the circuit = "Z_T =R + j(X_L - X_c) = 1000 +j(75-2652.52)"


"Z_T = 1000 + j2577.12 \space\space\varOmega"


Current through the circuit = "V\/Z_T = 106.07\/(1000-j2577.12) = 0.01388 +j0.03577"

There for current "I = 0.03837\angle68.79\degree"


We obsereve the rms value = "38.37 mA"


Maximum value of current = "rms value \times\sqrt{2} = 38.37\times\sqrt{2} = 54.26 m A"


The current can be written in sinusoidal form as "54.26Sin (377t+68.79\degree) m A"